# The Dual of Quantifier Elimination: Boolean Elimination over $\mathbb{C}$ and $\mathbb{R}$

### v2

@@ Abstract.
We show that every finite Boolean combination of polynomial equalities and inequalities in $\mathbb{C}^{n}$ admits two uniform normal forms: an $\exists\forall$ form and a $\forall\exists$ form, each using a single polynomial equation. Both forms have one existentially quantified variable and one universally quantified variable; regardless of the complexity of the original formula, no longer quantifier blocks are needed. The constructions are explicit and have linear degree bounds. Optimality results demonstrate that no purely existential or universal normal form is possible over $\mathbb{C}$.
Over $\mathbb{R}$, similar normal forms exist, including a singly-quantified $\exists$ form for Boolean combinations of equations and inequations, and $\exists^{d}$ and $\forall\exists$ forms for Boolean combinations involving order inequalities. Prior results establish the existence of a $\exists$ normal form for $\mathbb{R}$ by other methods. Finally, similar forms exist over $\mathbb{Q}$ as well.
These results may be viewed as a dual to classical quantifier elimination: instead of removing quantifiers at the cost of increased Boolean complexity, they remove Boolean structure at the cost of a short, fixed quantifier prefix.
@@ 1. Introduction
$\varphi(x)$ $\Longleftrightarrow(\exists a\in\mathbb{C})(\forall b\in\mathbb{C})\,p(a,b,x)=0$
$\varphi(x)$ $\Longleftrightarrow(\forall a\in\mathbb{C})(\exists b\in\mathbb{C})\,q(a,b,x)=0$
@@
Even more complicated formulas admit normal forms with only $\exists$ and one $\forall$, in either order. Furthermore the degrees of the resulting polynomials are linearly bounded by the length of the original formula in conjunctive or disjunctive normal form.
In logical terms, this shows that any Boolean combination of quantifier-free formulas can be expressed as a $\Delta_{2}$ formula with one atomic predicate.
The constructions below apply not only to polynomial terms but to any term language which includes polynomial combinations. For example, if $x$ and $y$ denote entire functions of a complex variable, then the resulting quantified equations describe entire functions as well, albeit of greater algebraic complexity.
These normal forms have application to model theory. For instance, in analyzing Zilber’s quasiminimality conjecture, our results show that it suffices to establish quasiminimality for sets defined by a single quantified equation, rather than sets defined by quantifications of arbitrary propositional combinations. More generally, having both normal forms provides flexibility for use inside larger formulas: the $\exists\forall$ form is suited to contexts where the innermost quantifier block is existential, and the $\forall\exists$ for contexts where the innermost block is universal.
@@
The paper is organized as follows: Section 2 establishes notation. Section 3 reviews context. Sections 4 and 5 construct the $\exists\forall$ and $\forall\exists$ polynomials for $\mathbb{C}$ explicitly, and provide bounds on their degrees. Section 6 shows that these quantifier patterns are optimal: there are simple formulas over $\mathbb{C}$ that cannot be expressed by purely existential or purely universal quantification of a single equation.
@@
Section 7 turns to other base fields, and shows that any Boolean combination of equations and inequations over $\mathbb{R}$ ($u\neq 0$) can be represented with a single $\exists$ and a single equality. Section 8 constructs the $\exists^{d}$ and $\forall\exists$ polynomials for Boolean combinations over $\mathbb{R}$ that may involve inequalities ($u>0$). Sections 9 and 10 explore prior results over $\mathbb{R}$ and forms over $\mathbb{Q}$. Section 11 concludes with directions for further research.
@@ 3. Context
We call the transformations of this paper Boolean elimination to highlight their duality with such quantifier elimination: instead of removing quantifiers at the cost of more propositional complexity, we remove propositional complexity at the cost of a short, fixed quantifier prefix.
@@ 5. Normalizing to $\forall\exists$ over $\mathbb{C}$
This $\varphi$ is equivalent to the $\forall\exists$ form below. Here we apply Lagrange interpolation to ensure that when $a$ equals $i$, only the term corresponding to the $i^{th}$ conjunct contributes to the polynomial.
@@ 6. Optimality: the quantifier prefixes cannot be shortened over $\mathbb{C}$
7. Extending the results on equations from $\mathbb{C}$ to $\mathbb{R}$
The constructions in theorems 4.1 and 5.1 above work not only over $\mathbb{C}$ but also in any infinite field of characteristic 0. The key properties used are the invertibility of elements, the infinitude of the field, and the finitude of roots for any nonzero polynomial. Thus the $\exists\forall$ and $\forall\exists$ normal forms hold over $\mathbb{R}$, $\mathbb{Q}$, $\mathbb{R}^{\text{alg}}$ and $\mathcal{O}$. Similarly, the argument of theorem 6.1 establishes that no purely universal normal form is possible over any of these fields.
A simple existential form is available for $\mathbb{R}$, or any other formally real field such as $\mathbb{Q}$, so long as we retain the restriction to formulas with equations and inequations. We switch notation for quantified variables from $a,b$ to $r,s$ to reflect the change from $\mathbb{C}$ to $\mathbb{R}$.
@@ 8. Normalizing to $\exists^{d}$ and $\forall\exists$ over $\mathbb{R}$
For inequalities in $\mathbb{R}$, it is typical to also consider order conditions such as $u>0$, and indeed Boolean combinations of such inequalities over $\mathbb{R}$ can be converted to both $\exists^{d}$ and $\forall\exists$ forms. We state theorems and constructions here analogous to the main theorem over $\mathbb{C}$, using the real-specific results in Lemma 3.1 with >> instead of $\neq$. We omit detailed proofs, especially since we will later cite other results establishing an $\exists$ form for these formulas.
@@
$\iff$ $\bigwedge_{i}\exists s;\prod_{j}t_{ij}\prod_{k}(1-s^{2}u_{ik})=0$
@@ 9. Single-quantifier $\exists$ forms over $\mathbb{R}$
As $r$ takes all possible values, graphs like the above sweep through the entire first quadrant.
@@ 11. Conclusion
  • With quantified input formulas, can every universally quantified formula be normalized to $\forall\exists$ form with a single equation, and every existentially quantified formula to $\exists\forall$ form?
  • What variants of these constructions are available when the Boolean combinations may be countable as in $\mathcal{L}_{\omega_{1},\omega}$ rather than finite, and the language may be expanded beyond polynomials to include exponentials or other univariate entire functions?

### v3

@@ Abstract.
We show that every finite Boolean combination of polynomial equalities and inequalities in $\mathbb{C}^{n}$ admits two uniform normal forms: an $\exists\forall$ form and a $\forall\exists$ form, each using a single polynomial equation. Both forms use only one existentially quantified variable and one universally quantified variable. Optimality results demonstrate that no purely existential or universal normal form is possible over $\mathbb{C}$.
These results extend to sets constructible from entire functions, and to quantifier-free formulas in functional languages over infinite fields of characteristic 0. A corollary shows Zilber’s conjecture on quasiminimality equivalent to its subcase quantifying a single equation.
Over $\mathbb{R}$ and $\mathbb{Q}$, similar results hold, including a singly-quantified $\exists$ form for Boolean combinations of equations and inequations, an $\exists$ form for $\mathbb{R}$ established by prior methods, and other results for order inequalities parallel to the forms over $\mathbb{C}$.
These results provide a dual to classical quantifier elimination: instead of removing quantifiers at the cost of increased Boolean complexity, they remove Boolean structure at the cost of a short, fixed quantifier prefix. The constructions have linear degree bounds and are explicit.
@@ 1. Introduction
$\varphi(x)$ $\Longleftrightarrow(\exists a\in\mathbb{C})(\forall b\in\mathbb{C})\,p(x,a,b)=0$
$\varphi(x)$ $\Longleftrightarrow(\forall a\in\mathbb{C})(\exists b\in\mathbb{C})\,q(x,a,b)=0$
@@
Even more complicated formulas admit normal forms with only $\exists$ and one $\forall$, in either order; having both normal forms provides flexibility for use inside larger formulas. Furthermore the degrees of the resulting polynomials are linearly bounded by the length of the original formula in conjunctive or disjunctive normal form.
The main result may also be generalized complex-geometrically or model-theoretically, with the same proofs.
Theorem 1.2.
[Reformulation in Complex Geometry] Let $\mathcal{A}$ be an algebra of entire functions on $\mathbb{C}^{n}$, closed under polynomial operations. Let $\mathcal{A}[a,b]$ denote the extension of this algebra by polynomials in two complex variables $a$ and $b$. Let $S\subset\mathbb{C}^{n}$ be any constructible set defined over $\mathcal{A}$, i.e. a finite Boolean combination of zero-sets $\{x:f(x)=0\}$ where $f\in\mathcal{A}$. Then $S$ can be represented as a locus of identical vanishing: there is a function $P\in\mathcal{A}[a,b]$ such that
$x\in S\Longleftrightarrow\exists a\in\mathbb{C}\text{ where }P(x,a,b)\text{ vanishes identically for all }b\in\mathbb{C}.$
Dually, $S$ can be represented as a locus of universal solvability: there is a function $Q\in\mathcal{A}[a,b]$ such that
[…]
The applies to several possible algebras, including the algebra of polynomials, the algebra of exponential polynomials, the algebra of entire functions of order $\rho$ (for some $\rho>0$), the algebra of entire functions of finite order, and the algebra of all entire functions.
Theorem 1.3.
[Reformulation in Model Theory] Let $\mathcal{L}$ be a functional language containing $1,+,-,\times$. Let $\mathcal{M}$ be an $\mathcal{L}$-structure satisfying the axioms for an infinite field of characteristic 0. Let $\varphi$ be any quantifier-free formula in $\mathcal{L}$ over the variables $x_{1},\ldots,x_{n}$. Then there are $\mathcal{L}$-terms $p,q$ formed from polynomial combinations of $1,a,b$ and the terms occurring in $\varphi$, such that
$\mathcal{M}\models\varphi(x)$ $\Longleftrightarrow(\exists a)(\forall b)\,p(x_{1},\ldots,x_{n},a,b)=0$
$\mathcal{M}\models\varphi(x)$ $\Longleftrightarrow(\forall a)(\exists b)\,q(x_{1},\ldots,x_{n},a,b)=0$
This applies to several possible languages, including the pure polynomial language, the language adding $\exp$, the language adding $1/\Gamma$, the language adding all univariate entire functions, and the language of all entire functions in any number of variables. In particular, we obtain an equivalence for Zilber’s conjecture on the quasiminimality of $\exp$ (on which see Kirby’s survey for more context).
Corollary 1.1.
Let $\mathcal{L}$ be the language $\{+,-,\times,\exp\}$, and let quantifiers $Q_{i}$ be either $\forall$ or $\exists.$ Then the following are equivalent:
  • For any $n$, any sequence of quantifiers $Q_{i}$, and any quantifier-free formula $\varphi(z,x_{1},\ldots,x_{n})$ in $\mathcal{L}$, the set
    $\{z\in C:Q_{1}x_{1}\ldots Q_{n}x_{n}\,\varphi(z,x_{1},\ldots,x_{n})\}\phantom{=0}$
    is either countable or cocountable.
  • For any $n$, any sequence of quantifiers $Q_{i}$, and any term $t(z,x_{1},\ldots,x_{n})$ in $\mathcal{L}$, the set
    $\{z\in C:Q_{1}x_{1}\ldots Q_{n}x_{n}\,t(z,x_{1},\ldots,x_{n})=0\}$
    is either countable or cocountable.
In other words, Zilber’s conjecture is equivalent to its subcase on the quasiminimality of sets defined by quantifying a single equation, rather than sets defined by quantifying an arbitrary formula.
@@
The rest of the paper is organized as follows: Section 2 establishes notation. Section 3 reviews context. Sections 4 and 5 construct the $\exists\forall$ and $\forall\exists$ polynomials for $\mathbb{C}$ explicitly, and provide bounds on their degrees. Section 6 shows that these quantifier patterns are optimal: there are simple formulas over $\mathbb{C}$ that cannot be expressed by purely existential or purely universal quantification of a single equation.
@@
Section 7 turns to other formally real fields, and shows that any Boolean combination of equations and inequations ($u\neq 0$) over them can be represented with a single $\exists$ and a single equality. Section 8 constructs the $\exists^{d}$ and $\forall\exists$ polynomials for Boolean combinations over $\mathbb{R}$ that may involve inequalities ($u>0$). Sections 9 and 10 explore prior results over $\mathbb{R}$ and forms over $\mathbb{Q}$. Section 11 concludes with directions for further research.
@@ 3. Context
We call the transformations of this paper Boolean elimination because they provide a dual to quantifier elimination: instead of removing quantifiers at the cost of more propositional complexity, we remove propositional complexity at the cost of a short, fixed quantifier prefix.
@@ 5. Normalizing to $\forall\exists$ over $\mathbb{C}$
We use $\mathbb{C}$ having characteristic 0 to assign a distinct integer parameter to each conjunct, and apply Lagrange interpolation to construct a polynomial whose $i^{th}$ term contributes only when $a$ equals $i$. Then $\varphi$ is equivalent to the $\forall\exists$ form below:
@@ 6. Optimality: the quantifier prefixes cannot be shortened over $\mathbb{C}$
7. Normalizing to $\exists$ over formally real fields
A simpler existential form, with only one new quantifier, is available for any other formally real field, so long as we retain the restriction to formulas with equations and inequations (but not order). This $\exists$ normal form holds for $\mathbb{R}$, $\mathbb{Q}$, and $\mathbb{R}^{\text{alg}}$.
This contrasts with the greater generality of theorems 4.1, 5.1, and 6.1, which work not only over $\mathbb{C}$ but also over any infinite field of characteristic 0. The key properties used there were the invertibility of elements, the infinitude of the field, and the finitude of roots for any nonzero polynomial. Here, by contrast, we use the defining property of a formally real field, namely that a sum of squares is zero iff all the squares are zero.
We switch notation for quantified variables from $a,b$ to $r,s$ to reflect the change from $\mathbb{C}$ to a field like $\mathbb{R}$.
@@ 8. Normalizing to $\exists^{d}$ and $\forall\exists$ over $\mathbb{R}$
For inequalities in $\mathbb{R}$, it is typical to also consider order conditions such as $u>0$. Indeed Boolean combinations polynomial equations and such inequalities over $\mathbb{R}$ can be converted to both $\exists^{d}$ and $\forall\exists$ forms. We state theorems here analogous to the main theorem over $\mathbb{C}$, using the algebraic encoding of order from Lemma 3.1. We omit detailed proofs, since the constructions and proofs are otherwise parallel to the results over $\mathbb{C}$, and we will later cite other results establishing a single-quantifier $\exists$ form for these formulas.
@@
$\iff$ $\bigwedge_{i}\exists s\,\prod_{j}t_{ij}\prod_{k}(1-s^{2}u_{ik})=0$
@@ 9. Single-quantifier $\exists$ forms over $\mathbb{R}$
As $r$ takes all possible values, graphs like the below sweep through the entire first quadrant.
@@
A contrast between Pecker’s results and ours is that our results, by using two new variables, require only one use of each term $t_{ij}$ or $u_{ik}$, and have linearly growing degrees. Pecker’s results, using only one new variable, require multiple uses of those terms, and have exponentially growing degrees.
@@ 11. Conclusion
  • Can every universally quantified formula be normalized to $\forall\exists$ form with a single equation, and can every existentially quantified formula be normalized to $\exists\forall$ form?
  • What variants of these constructions are available over $\mathbb{C}$ when the Boolean combinations may be countable as in $\mathcal{L}_{\omega_{1},\omega}$ rather than finite, and the language may be expanded beyond polynomials to include exponentials or other univariate entire functions?